Light is incident on a diffraction grating that has 300 lines /mm. A diffraction pattern is observed. The angle between the central maximum and second order maximum was measured to be 250 . Calculate the wavelength of the light.

1
Expert's answer
2022-06-07T08:11:55-0400

"d\sin\varphi=k\lambda,"

"\lambda=\frac{d\sin\varphi}k=704.4~nm."